Star Videos, Inc., produces short musical videos for sale to retail outlets. The company’s balance sheet accounts as of January 1 are given below.

 

Star Videos, Inc.
Balance Sheet
January 1
Assets




Cash


$89,000
Accounts receivable



104,400
Inventories:




Raw materials (film, costumes)$20,200


Videos in process
59,000


Finished videos awaiting sale
82,400

161,600
Prepaid insurance



12,350
Studio and equipment (net)



618,000
Total assets


$985,350
Liabilities and Stockholders’ Equity




Accounts payable


$180,000
Retained earnings



805,350
Total liabilities and stockholders’ equity


$985,350

 

 

Because the videos differ in length and in complexity of production, the company uses a job-order costing system to determine the cost of each video produced. Studio (manufacturing) overhead is charged to videos on the basis of camera-hours of activity. The company’s predetermined overhead rate for the year ($40 per camera-hour) is based on a cost formula that estimated $280,000 in manufacturing overhead for an estimated allocation base of 7,000 camera-hours. Any underapplied or overapplied overhead is closed to cost of goods sold. The following transactions were recorded for the year:

 

  1. Film, costumes, and similar raw materials purchased on account, $194,500.

  2. Film, costumes, and other raw materials issued to production, $210,000 (85% of this material was considered direct to the videos in production, and the other 15% was considered indirect).

  3. Utility costs incurred (on account) in the production studio, $95,400.

  4. Depreciation recorded on the studio, cameras, and other equipment, $88,400. Three-fourths of this depreciation related to actual production of the videos, and the remainder related to equipment used in marketing and administration.

  5. Advertising expense incurred (on account), $139,500.

  6. Salaries and wages paid in cash as follows:

 




Direct labor (actors and directors)$95,400
Indirect labor (carpenters to build sets, costume designers, and so forth)$82,000
Administrative salaries$97,400

 

  1. Prepaid insurance expired during the year, $10,950 (70% related to production of videos, and 30% related to marketing and administrative activities).

  2. Miscellaneous marketing and administrative expenses incurred (on account), $13,650.

  3. Studio (manufacturing) overhead was applied to videos in production. The company recorded 7,250 camera-hours of activity during the year.

  4. Videos that cost $504,000 to produce according to their job cost sheets were transferred to the finished videos warehouse to await sale and shipment.

  5. Sales for the year totaled $988,000 and were all on account. 

  6. The total cost to produce the videos that were sold according to their job cost sheets was $544,730.

  7. Collections from customers during the year totaled $938,000.

  8. Payments to suppliers on account during the year, $528,000.

  9. Underapplied or overapplied overhead $__?__.

 

Required:

1. Prepare a transaction analysis that records all of the above transactions.

2. Prepare a schedule of cost of goods manufactured for the year.

3. Prepare a schedule of cost of goods sold for the year.

4. Prepare an income statement for the year.
Answer #1

1.

Star Videos, Inc.
Transaction Analysis
For the Year Ended December 31
Cash Accounts Receivable Raw Materials Videos in Process Finished Videos Manufacturing Overhead Prepaid Insurance Studio & Equipment (net) = Accounts Payable Retained Earnings
Beginning balance @ 1/1 89000 104400 20200 59000 82400 0 12350 618000 = 180000 805350
a) Raw material purchases 194500 = 194500
b) Raw materials used -210000 178500 31500 =
c) Utility costs 95400 = 95400
d) Depreciation charges 66300 -88400 = -22100
e) Advertising = 139500 -139500
f) Salaries & wages -274800 95400 82000 = -97400
g) Prepaid insurance 7665 -10950 = -3285
h) Miscellaneous marketing = 13650 -13650
i) Applied overhead 290000 -290000 =
j) Transfer completed videos to finished goods -504000 504000 =
k) Sales 988000 = 988000
l) Transfer finished goods to cost of goods sold -544730 = -544730
m) Cash collections from customers 938000 -938000 =
n) Payment to suppliers -528000 = -528000
o) Overapplied overhead 7135 = 7135
Ending balances @12/31 224200 154400 4700 118900 41670 0 1400 529600 = 95050 979820

2.

Star Videos, Inc.
Schedule of Cost of Goods Manufactured
For the Year Ended December 31
Beginning raw materials inventory $ 20200
Add: Raw material purchases 194500
Raw materials available for use 214700
Less: Ending raw materials inventory 4700
Raw materials used 210000
Less: Indirect materials 31500 178500
Direct labor 95400
Manufacturing overhead 290000
Total manufacturing costs 563900
Add: Beginning videos in process inventory 59000
Total videos in process inventory 622900
Less: Ending videos in process inventory 118900
Cost of goods manufactured 504000

3.

Star Videos, Inc.
Schedule of Cost of Goods Sold
For the Year Ended December 31
Cost of goods manufactured 504000
Add: Beginning finished goods inventory 82400
Cost of goods available for sale 586400
Less: Ending finished goods inventory 41670
Cost of goods sold 544730
Less: Overapplied overhead 7135
Adjusted cost of goods sold $ 537595

4.

Star Videos, Inc.
Income Statement
For the Year Ended December 31
Sales revenue 988000
Cost of goods sold 537595
Gross profit 450405
Operating expenses 275935
Net income $ 174470